# SyncDiff: Synchronized Motion Diffusion for Multi-Body Human-Object Interaction Synthesis

### v2

@@ Abstract
Synthesizing realistic human-object interaction motions is a critical problem in VR/AR and human animation. Unlike the commonly studied scenarios involving a single human or hand interacting with one object, we address a more generic multi-body setting with arbitrary numbers of humans, hands, and objects. This complexity introduces significant challenges in synchronizing motions due to the high correlations and mutual influences among bodies. To address these challenges, we introduce SyncDiff, a novel method for multi-body interaction synthesis using a synchronized motion diffusion strategy. SyncDiff employs a single diffusion model to capture the joint distribution of multi-body motions. To enhance motion fidelity, we propose a frequency-domain motion decomposition scheme. Additionally, we introduce a new set of alignment scores to emphasize the synchronization of different body motions. SyncDiff jointly optimizes both data sample likelihood and alignment likelihood through an explicit synchronization strategy. Extensive experiments across four datasets with various multi-body configurations demonstrate the superiority of SyncDiff over existing state-of-the-art motion synthesis methods.
SyncDiff is a unified framework synthesizing synchronized multi-body interaction motions with any number of hands, humans, and rigid objects. The key innovations in SyncDiff are two novel multi-body motion synchronization mechanisms, namely the alignment scores and corresponding alignment loss function $\mathcal{L}_{\text{align}}$, and explicit synchronization strategy in inference. These mechanisms can help SyncDiff synthesize realistic, coordinated, and physically plausible motions in various interaction scenarios.
22footnotetext: Corresponding author.
@@ 1 Introduction
In everyday life, humans frequently manipulate objects to complete tasks, often using both hands or collaborating with others. For instance, an individual might use both hands to set a fallen chair upright, hold a brush in one hand to scrub a bowl held in the other, or work with another person to lift and position a heavy object. These activities are examples of multi-body human-object interactions (HOI) [40, 13, 39], where “body” may refer to objects, hands, or humans. Synthesizing such interactions is a prominent research area with significant applications in VR/AR, human animation, and robot learning [88].
@@
The primary challenge in multi-body HOI synthesis is capturing the complex joint distribution of body motions and ensuring that the individual body motions are not only synchronized but also mutually aligned with meaningful interaction semantics. This challenge intensifies as the number of bodies increases with high-order motion relationships. Existing works usually focus on specific configurations of the bodies with limited body numbers such as a single hand interacting with an object [88], a person engaging with an object [40, 78], or two hands manipulating a single item [68, 13]. These works introduce configuration-specific priors and are restricted to particular setups, leading to a strong demand for a generic method that could handle the series of multi-body HOI synthesis problems in a unified manner without restrictions for body numbers.
@@
To develop a generic solution, we draw two key insights. First, there exists a straightforward method of treating the motions of all individual bodies as a high-dimensional representation and using a single diffusion model to depict its distribution [13]. However, this diffusion model only estimates the data sample scores, which only implicitly depicts the correlations across individual body motions and is usually insufficient to guarantee the alignment between individual body motions with limited data amount. To solve this, we need to explicitly model a new set of alignment scores, which guides the model to synthesize synchronized motions. Second, once we estimate the sample and alignment scores from diffusion models, in inference, we can jointly optimize sample and alignment likelihoods by approaching multi-body HOI synthesis as a motion synchronization problem defined on a graphical model, whose nodes represent single bodies’ individual motions, and edges describe the relative motions between pairwise bodies.
To realize the ideas above, we design Synchronized Motion Diffusion (SyncDiff) for generic multi-body HOI synthesis. SyncDiff is a diffusion model defined in the graphical model above, which simultaneously estimates the data sample scores and alignment scores by operating on a redundant motion representation consisting of all individual motions and relative motions. In particular, we propose frequency-domain motion decomposition to better estimate sample scores, which factorizes all individual and relative motions into different frequency components and supervises high-frequency components generated in the frequency domain independently. This approach prevents high-frequency, small-amplitude vibrations from being overshadowed by low-frequency, large-scale movements. For better synchronization, we design a set of alignment scores, which measures the direction where computed relative motion between two bodies’ synthesized individual motions approaches the synthesized relative motion on the edge connecting them. At inference time, by leveraging both sample and alignment scores, we discover a simple explicit synchronization strategy, which is equivalent to maximum likelihood inference with a theoretical justification.
Integrating these motion synchronization and decomposition techniques, SyncDiff consistently improves multi-body interaction coordination and fine-grained motion fidelity. Experiments on four datasets involving different numbers of hands, humans, and objects demonstrate the strong generalizability of our method over many state-of-the-art methods in every specific setting.
In summary, we propose a newly adapted diffusion model(SyncDiff) defined on graphical models to synthesize synchronized motions. It incorporates a frequency-domain decomposition to capture fine-grained periodic motions. We also provide theoretical derivation for two synchronization mechanisms, namely the alignment scores and explicit synchronization in inference.
2 Related Work
@@ 2.1 Human-object Interaction Motion Synthesis
Synthesizing interactive motions between humans and objects has been an emerging research field supported by extensive HOI datasets  [2, 3, 5, 20, 24, 37, 45, 76, 94, 36, 98, 46, 99]. A branch of works [25, 54, 86, 87, 109] aims at synthesizing human bodies interacting with objects, while others [61, 21, 50, qin2022dexmvimitationlearningdexterous, 88, 112, 103, 4] focus on hand-object interactions for learning more dexterous manipulation behaviors. For synthesizing human-object interaction motions, early works [108, 22, 107, 77] propose to leverage conditional variational auto-encoders [35, 71] to model distributions of human-object interaction motions and enable generalization ability, meanwhile using CLIP [60] to encode language features for text-based control. With the tremendous progress of diffusion models [28], diffusion-based methods [78, 40, 39, 92, 56, 90, 19, 89] have been widely proposed for superior generation qualities. Focusing on multi-person and object collaboration synthesis, recent approaches [67, 42, 56, 14] enhance feature exchange among different entities in diffusion steps for better cross-entity synchronization. For synthesizing hand-object interaction motions, one solution is to train dexterous hand agents [10, 7, 101, 11, 102, 83] using physical simulations [82, 49] and reinforcement learning [23, 66]. Another solution that is similar to ours is fully data-driven. Early methodologies [43, 44, 93, 100, 110, 113] apply representation techniques such as contact map or reference grasp, which better model the contact between hand joints or hand surface points and the relevant local regions of the object. Several recent works [12, 68, 103, 4, 106] have shifted the focus to multi-hand collaboration, while few works have addressed interactions involving more than one object. Compared to the above methods, our framework can handle both human-object and hand-object interactions, allowing any number of bodies in the scene, without the need for any delicately designed structures like contact guidance.
@@ 2.2 Injecting External Knowledge into Diffusion Models
Benefitting from high generation qualities, diffusion models [28, 72] are widely adopted in versatile synthesis tasks regarding images [63, 104], videos [30, 29], and motions [78, 105]. To further induce generation results to satisfy specific demands or constraints, injecting additional knowledge (e.g., expected image styles, human-object contact constraints) into diffusion processes is an emerging methodology in recent studies [92, 39, 56, 19]. For this purpose, one paradigm is explicitly improving intermediate denoising results in inference steps using linear fusions with conditions [48, 9, 78], optimizations with differential energy functions [17, 39, 56, 96], or guidance from learnable networks [92, 97, 32]. Another paradigm is to design additional diffusion branches with novel conditions [27, 111, 91] or representations [19] that comprise the knowledge, encompassing the challenge of fusing them with existing branches. Classifier-free guidance [27, 111] uses a linear combination of results from different branches in inference steps, Xu et. al. [91] propose blending and learnable merging strategies, while CG-HOI [19] presents additional cross-attention modules. Our method follows the second paradigm with relative representation designs and transforms the problem into a graphical model, optimizing it with an explicit synchronization strategy.
@@ 2.3 Motion Synchronization
Motion Synchronization aims at coordinating the motions of different bodies in a multi-body system to satisfy specific inter-body constraints. As technical bases, for specific parameterized closed-form constraints, the closed-form solutions in Euler angles, axis-angle, quaternions, and rotation matrices are derived in traditional approaches [8, 34, 53, 6, 65, 31, 114, 1]. To achieve global attitude synchronization under mechanical constraints in $SE(3)$, several works adopt gradient flow [51, 52], lifting method [80], and matrix decomposition [81] techniques. In the trend of graph-based multi-body modeling, existing approaches present coordination control based on general undirected [62] and directed [84, 47, 85] graphs, while some works [26, 79] further extend to nonlinear configurations and dynamic topologies. These works do not adopt learning-based methods; by contrast, our method uses a generative model to enhance consistency, formulating data-driven motion synchronization.
@@ 2.4 Modeling Motions in Frequency Domain
3 Problem Formulation and Challenges
Considering an interaction scenario comprising $n$ articulated skeletons $h_{1},h_{2},\dots,h_{n}$ and $m$ rigid bodies $o_{1},o_{2},\dots,o_{m}$, where articulated skeletons are sets of joints allowing the skeletons’ motion to be fully reconstructed based on joints’ positions and intrinsic shape parameters. Our task is to synthesize their motions with known action labels, object categories, object models, and shape parameters $\beta_{i\in[1,n]}$ of each skeleton. The skeletons are MANO [64] hands or SMPL-X [55] humans in our experiments, and we would like to highlight that our task releases the needs of either predefined object motions [40] or hand-object key frames [13] that are used in existing works.
Compared to previous single-object interaction motion synthesis problems [13, 40, 68], two new challenges lie in synthesizing multi-skeleton multi-object interactions. Firstly, governed by abundant topologies, geometries, and versatile functionalities of objects, collaborations among multiple objects are more diverse and complex than those of one hand/human and one object [112], which typically follow constant grasping poses and can easily benefit from learnable priors (e.g., contact maps [44, 3, 38, 93], affordance maps [33, 50, 95, 56]). Secondly, the increase in bodies and objects puts forward higher demands for multi-entity motion synchronization. Slight temporal misalignment between any two entities could dramatically lead to unrealistic and failing cooperation behaviors.
@@ 4 Method
As mentioned in Section 1, in SyncDiff, representation factorization (Section 4.1) involves two aspects. On one hand, in the graphical model, we need individual motions for single bodies on nodes, and relative motions for pairwise bodies on edges. On the other hand, individual/relative representations are both decomposed into low- and high-frequency components. For our synchronization mechanisms, we will elaborate on both the implicit joint modeling (Section 4.2) and the explicit synchronization strategy during inference (Section 4.3). In order to produce hand mesh or human body mesh from predicted joints, the post processing algorithm of mesh construction is introduced (Section 4.4).
@@
Figure 1: Overview of SyncDiff. The light blue box shows the inference process with explicit synchronization performed every $t_{0}$ step. For denoising steps irrelevant to explicit synchronization, the noise level is set to $\sigma_{t}$. For the calculation from $\hat{x}$ to $\hat{\mu}$, no noise is added. For the denoising process that involves explicit synchronization, the noise level is $\sigma_{t}^{{}^{\prime}}$. Please refer to Section 4.3 for more details. The light green box illustrates denoising model architecture (Section 4.2.1).
4.1 Factorized Motion Representations
@@
1. Individual and Relative representations. For a single body $\textbf{b}=h_{i\in[1,n]}$ or $\textbf{b}=o_{j\in[1,m]}$, we represent its motion as $x_{u,\textbf{b}}$ in the world coordinate system. Here $x_{u,h_{i}}\in\mathbb{R}^{T\times D\times 3}$, where $D$ is the number of its joints, represents 3D joint positions. $x_{u,o_{j}}=[t_{j},q_{j}]\in\mathbb{R}^{T\times 7}$, where $t_{j}\in\mathbb{R}^{T\times 3}$ is its translation, and $q_{j}\in\mathbb{R}^{T\times 4}$ is the quaternion representing its orientation.
@@
For two bodies $\textbf{b}_{1}$ and $\textbf{b}_{2}$, we use $x_{b,\textbf{b}_{2}\rightarrow\textbf{b}_{1}}$ to denote the relative motion of $\textbf{b}_{2}$ in $\textbf{b}_{1}$’s coordinate system. We require that $\textbf{b}_{1}$ is a rigid body, and omit relative representation between articulated skeletons. If $\textbf{b}_{2}$ is also a rigid body, then $x_{b,\textbf{b}_{2}\rightarrow\textbf{b}_{1}}=[t_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}},q_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}]$, where $t_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}=q_{\textbf{b}_{1}}^{-1}(t_{\textbf{b}_{2}}-t_{\textbf{b}_{1}})$, and $q_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}=q_{\textbf{b}_{1}}^{-1}q_{\textbf{b}_{2}}$. If $\textbf{b}_{2}$ is an articulated skeleton, the position of one of its joints $p$ under individual representation will become $p^{{}^{\prime}}$ in $\textbf{b}_{1}$’s coordinate system, where $p^{{}^{\prime}}=q_{\textbf{b}_{1}}^{-1}(p-t_{\textbf{b}_{1}})$.
@@
After deriving all individual and relative motion representations, we concatenate all of them into $x=[x_{u,o_{1}},x_{u,o_{2}},\dots,x_{u,o_{m}},x_{u,h_{1}},x_{u,h_{2}},\dots,x_{u,h_{n}},\\ x_{b,o_{2}\rightarrow o_{1}},x_{b,o_{3}\rightarrow o_{1}},\dots,x_{b,o_{m}\rightarrow o_{m-1}},x_{b,h_{1}\rightarrow o_{1}},x_{b,h_{1}\rightarrow o_{2}},\\ \dots,x_{b,h_{n}\rightarrow o_{m}}]\in\mathbb{R}^{T\times D_{\text{sum}}}$, where $D_{\text{sum}}=7m+3Dn+7{m\choose 2}+3Dmn$. Although $x$ is composed of individual and relative representations, for final synthesized motions of bodies, they only involve individual motions, and those relative ones serve merely as auxiliary components.

### v3

@@ Abstract
Synthesizing realistic human-object interaction motions is a critical problem in VR/AR and human animation. Unlike the commonly studied scenarios involving a single human or hand interacting with one object, we address a more generic multi-body setting with arbitrary numbers of humans, hands, and objects. The high correlations and mutual influences among bodies leads to two major challenges, for which we propose solutions. First, to satisfy the high demands for synchronization of different body motions, we mathematically derive a new set of alignment scores during the training process, and use maximum likelihood sampling on a dynamic graphical model for explicit synchronization during inference. Second, the high-frequency interactions between objects are often overshadowed by the large-scale low-frequency movements. To address this, we introduce frequency decomposition and explicitly represent high-frequency components in the frequency domain. Extensive experiments across five datasets with various multi-body configurations demonstrate the superiority of SyncDiff over existing state-of-the-art motion synthesis methods.
Figure 1: SyncDiff is a unified framework synthesizing synchronized multi-body interaction motions with any number of hands, humans, and rigid objects. In SyncDiff, we introduce two novel multi-body motion synchronization mechanisms, namely the alignment scores for training and explicit synchronization strategy in inference. With these mechanisms, the synthesized results can effectively prevent interpenetration, contact loss, or asynchronous human-object interactions in various scenarios, as shown in the above figure.
@@ 1 Introduction
In daily life, humans frequently manipulate objects to complete tasks, often using both hands or collaborating with others. For instance, an individual might use both hands to set up a fallen chair, hold a brush in one hand to scrub a bowl held in the other, or work with another person to lift and position a heavy object. These are examples of multi-body human-object interactions (HOI) [42, 13, 41], where “body” can refer to objects, hands, or humans. Synthesizing such interactions has prominent applications in VR/AR, human animation, and robot learning [96].
@@
The primary challenge in multi-body HOI synthesis is capturing the complex joint distribution of body motions and ensuring that the individual body motions are not only synchronized but also mutually aligned with meaningful interaction semantics. This challenge intensifies as the number of bodies increases with high-order motion relationships. Existing works only focus on some specific configurations with limited body numbers, such as a single hand interacting with an object [96], a person engaging with an object [42, 84], or two hands manipulating a single item [73, 13], where motion trajectories are often simple and without too many semantics. These works introduce configuration-specific priors and are restricted to particular setups, leading to a strong demand for a generic method that could handle the series of multi-body HOI synthesis problems in a unified manner without restrictions for body numbers, and with more complex motion distributions.
@@
A straightforward method to promote synchronization is treating the motions of all individual bodies as a high-dimensional representation and using a single diffusion model to depict its distribution [13]. However, this diffusion model only estimates data sample scores, and only implicitly depicts the correlations across individual motions. To further promote the alignment between individual body motions with limited data amount, analogous to the data sample scores guiding data denoising and reconstruction, we need a set of alignment scores to promote motion synchronization. Second, with estimated sample and alignment scores, in inference, we can jointly optimize sample and alignment likelihoods by formulating multi-body HOI synthesis as a motion synchronization problem within a graphical model, where nodes represent individual body motions and edges capture the relative motions between body pairs.
In practice, we often observe high-frequency, small-amplitude vibrations with semantics be overshadowed by low-frequency, large-scale movements (e.g., A brush only contacts the teapot, without relative frictions between them). To solve this, we propose frequency decomposition, which factorizes all individual and relative motions into low and high frequency components, and supervises high-frequency ones in the frequency domain independently.
To realize the ideas above, we design Synchronized Motion Diffusion (SyncDiff) for generic multi-body HOI synthesis. SyncDiff is a diffusion model defined on the graphical model above, operating on a high-order motion representation consisting of all individual motions and relative motions, which is the first unified framework for multi-body HOI synthesis of any body count. Our detailed contributions include: 1) For better synchronization, we derive a set of alignment scores and the corresponding loss term for training. 2) At inference time, by leveraging both sample and alignment scores, we discover a simple explicit synchronization strategy, which is equivalent to maximum likelihood inference with a mathematical justification. 3) We decompose all individual and relative motions based on frequency, to better model high-frequency components with semantics. 4) Extensive experiments on five datasets demonstrate the superior human-object contact quality of our method, and the semantic recognition accuracy rate surpasses the sota methods by an average of over 15%.
2 Related Works
@@ 2.1 Human-object Interaction Motion Synthesis
Synthesizing interactive motions between humans and objects has been an emerging research field supported by extensive HOI datasets  [3, 4, 6, 21, 26, 40, 47, 81, 105, 39, 108, 48, 109, 100, 2, 119]. A branch of works [27, 56, 93, 94, 121] aims at synthesizing human bodies interacting with objects, while others [65, 23, 52, 63, 96, 124, 113, 5, 34] focus on hand-object interactions for learning more dexterous manipulation behaviors. For synthesizing human-object interaction motions, early works [120, 24, 117, 82, 62] propose to leverage conditional variational auto-encoders [38, 76] to model distributions of human-object interaction motions and enable generalization ability, meanwhile using CLIP [64] to encode language features for text-based control. With the tremendous progress of diffusion models [30], diffusion-based methods [84, 42, 41, 102, 58, 98, 20, 97, 14, 95] have been widely proposed for superior generation qualities. Focusing on multi-person and object collaboration synthesis, recent approaches [71, 44, 58, 15, 83, 101, 118] enhance feature exchange among different entities in diffusion steps for better cross-entity synchronization. For synthesizing hand-object interaction motions, one solution is to train dexterous hand agents [11, 8, 111, 112, 90, 103] using physical simulations [89, 51] and reinforcement learning [25, 70]. Another solution that is similar to ours is fully data-driven. Early methodologies [45, 46, 104, 110, 122, 125] apply representation techniques such as contact map or reference grasp, which better model the contact between hand joints or hand surface points and the relevant local regions of the object. Several recent works [12, 73, 113, 5, 116] have shifted the focus to multi-hand collaboration, while few works have addressed interactions involving more than one object. Compared to the above methods, our framework can handle both human-object and hand-object interactions, allowing any number of bodies in the scene, without the need for any delicately designed structures like contact guidance.
@@ 2.2 Injecting External Knowledge into Diffusion Models
Benefitting from high generation qualities, diffusion models [30, 77] are widely adopted in versatile synthesis tasks regarding images [67, 114], videos [32, 31], and motions [84, 115]. To further induce generation results to satisfy specific demands or constraints, injecting additional knowledge (e.g., expected image styles, human-object contact constraints) into diffusion processes is an emerging methodology in recent studies [102, 41, 58, 20]. For this purpose, one paradigm is explicitly improving intermediate denoising results in inference steps using linear fusions/imputations with conditions [50, 10, 84, 37], optimizations with differential energy functions [18, 41, 58, 106], or guidance from learnable networks [102, 107, 35, 22, 88]. Another paradigm is to design additional diffusion branches with novel conditions [29, 123, 99] or representations [20] that comprise the knowledge, encompassing the challenge of fusing them with existing branches. Classifier-free guidance [29, 123] uses a linear combination of results from different branches in inference steps, Xu et al. [99] propose blending and learnable merging strategies, while CG-HOI [20] presents additional cross-attention modules. Our method follows the second paradigm with relative representation designs and transforms the problem into a graphical model, optimizing it with an explicit synchronization strategy.
@@ 2.3 Motion Synchronization
Motion Synchronization aims at coordinating motions of different bodies in a multi-body system to satisfy specific inter-body constraints. As technical bases, for specific parameterized closed-form constraints, the solutions in Euler angles, axis-angle, quaternions, and rotation matrices are derived in traditional approaches [9, 36, 55, 7, 69, 33, 127, 1]. To achieve global attitude synchronization under mechanical constraints in $SE(3)$, several works adopt gradient flow [53, 54], lifting method [86], and matrix decomposition [87] techniques. In the trend of graph-based multi-body modeling, existing approaches present coordination control based on general undirected [66] and directed [91, 49, 92] graphs, while some works [28, 85] further extend to nonlinear configurations and dynamic topologies. These works do not adopt learning-based methods; by contrast, our method uses a generative model to enhance consistency, formulating data-driven motion synchronization.
@@ 3 Method
In this section, we first formulate the problem (Section  3.1), and describe the motion representation we use (Section  3.2). Then we introduce our algorithm of frequency decomposition (Section  3.3) and model architecture (Section  3.4) in detail. Finally, in Section  3.5 and  3.6, our two synchronization mechanisms will be elaborated.
@@
Figure 2: Overview of SyncDiff. The light blue boxes show the inference process with explicit synchronization steps performed every $s$ step. For denoising steps irrelevant to explicit synchronization (those marked as “$(s-1)$ times”), the noise level is set to $\sigma_{t}$. For the calculation from $\hat{x}$ to $\hat{\mu}$, no noise is added. For the calculation of $\hat{x}_{t-1}$ based on $\hat{x}_{t}$ and $\hat{\mu}_{t}$ in explicit synchronization step, the noise level is $\sigma_{t}^{{}^{\prime}}$. Please refer to Section 3.6 for more details. The light green box illustrates the architecture of denoising backbone (Section 3.4).
3.1 Problem Definition
Consider an interaction scenario comprising $n$ articulated skeletons $h_{1},h_{2},\dots,h_{n}$ and $m$ rigid bodies $o_{1},o_{2},\dots,o_{m}$, which we uniformly define as “body”. In our experiments, these skeletons are MANO [68] hands or SMPL-X [57] humans, whose motions can be reconstructed with joint information and intrinsic shape parameter $\beta$. Details can be found in Section  C.2. Our task is to synthesize body motions with known action labels, object categories and geometries, and shape parameters $\beta_{i\in[1,n]}$ of each skeleton. We highlight that our task releases the needs of either predefined object motions [42] or HOI key frames [13] that are used in other existing works.
3.2 Motion Representation
As mentioned in Section 1, in the graphical model, we need individual motions for single bodies on nodes, and relative motions for pairwise bodies on edges.
@@ 3.2 Motion Representation
For a single body $\textbf{b}=h_{i\in[1,n]}$ or $\textbf{b}=o_{j\in[1,m]}$, we represent its motion as $x_{\textbf{b}}$ in the world coordinate system. Here $x_{h_{i}}\in\mathbb{R}^{N\times 3D}$ represents 3D joint positions, where $N$ is number of frames, and $D$ is number of joints. $x_{o_{j}}=[\textbf{t}_{o_{j}},\textbf{q}_{o_{j}}]\in\mathbb{R}^{N\times 7}$, where $\textbf{t}_{o_{j}}\in\mathbb{R}^{N\times 3}$ is its translation, and $\textbf{q}_{o_{j}}\in\mathbb{R}^{N\times 4}$ is the quaternion for its orientation.
@@
For two bodies $\textbf{b}_{1}$ and $\textbf{b}_{2}$, we use $x_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}$ to denote the relative motion of $\textbf{b}_{2}$ in $\textbf{b}_{1}$’s coordinate system. We require that $\textbf{b}_{1}$ is a rigid body, and omit relative representation between articulated skeletons. If $\textbf{b}_{2}$ is also a rigid body, then
$x_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}=[\textbf{t}_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}},\textbf{q}_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}]=[\textbf{q}_{\textbf{b}_{1}}^{-1}(\textbf{t}_{\textbf{b}_{2}}-\textbf{t}_{\textbf{b}_{1}}),\textbf{q}_{\textbf{b}_{1}}^{-1}\textbf{q}_{\textbf{b}_{2}}].$ (1)
If $\textbf{b}_{2}$ is an articulated skeleton, the position of one of its joints p under individual representation will become $\textbf{p}^{{}^{\prime}}$ in $\textbf{b}_{1}$’s coordinate system, where
$\textbf{p}^{{}^{\prime}}=\textbf{q}_{\textbf{b}_{1}}^{-1}(\textbf{p}-\textbf{t}_{\textbf{b}_{1}}).$ (2)
After deriving all individual and relative motion representations, we concatenate all of them together into a high-order representation $x$, including $\{x_{o_{j\in[1,m]}}\}$, $\{x_{h_{i\in[1,n]}}\}$, $\{x_{o_{j_{2}}\rightarrow o_{j_{1}}}\mid j_{1},j_{2}\in[1,m],j_{1}\neq j_{2}\}$, $\{x_{h_{i}\rightarrow o_{j}}\mid i\in[1,n],j\in[1,m]\}$. Thus, $x\in\mathbb{R}^{N\times D_{\text{sum}}}$, where
$D_{\text{sum}}$ $=7m+3Dn+7m(m-1)+3Dmn$
$=7m^{2}+3Dn(m+1).$ (3)
@@
Although $x$ is composed of individual and relative representations, the final synthesized motions only involve individual motions, and those relative ones serve merely as auxiliary components.
3.3 Frequency Decomposition
As body number increases, action semantics become more dependent on subtle, high-frequency interactions between bodies (e.g., periodic frictions between the teapot and brush during the action of brushing). Unfortunately, using existing works to denoise based on time-domain trajectories often results in these high-frequency components being overshadowed by common low-frequency, large-scale patterns (e.g., objects only move and contact each other, without relative interactions). To better model these high-frequency components, inspired by GID [43], we explicitly represent them in the frequency domain and supervise them independently in the loss terms described in Section  3.5.